im having trouble understanding how they get get the answer i was hoping maybe you can help please


F=9C+32 find F if C =60 i dont see how they are getting F=180 i got
5
everything after that.
thankyou, Nicole

Answer by checkley75(3666)

F=9C/5+32
F=9*60/5+32
F=540/5+32
F=108+32
F=140
I ALSO DON'T SEE HOW ANYONE GETS F=180.